A capacitor with very large capacitance is in series with another capacitor with very small capacitance.
What is the equivalent capacitance of the combination?
(a)Slightly greater than the capacitance of the large capacitor
(b) Slightly less than the capacitor of the large capacitor
(c) Slightly greater than the capacitance of the small capacitor
(d) Slightly less than the capacitance of the small capacitor
1
Expert's answer
2021-02-17T11:10:17-0500

The total capacity is given by "\frac{1}{C} = \frac{1}{C_1}+\frac{1}{C_2}" and thus is smaller than both "C_1, C_2" as "1/C \geq \max(1/C_1,1/C_2), C \leq \min (C_1, C_2)" . Therefore the answer is d.